OpenStudy (anonymous):

How to solve this fractional equation x^2-2x+1/x-2=0?

Parth (parthkohli):

First, divide \(x - 2\) from \(x^2 - 2x + 1\), then solve that equation.

OpenStudy (anonymous):

How?

OpenStudy (helder_edwin):

can u divide by zero?

OpenStudy (helder_edwin):

well u can't. so when you have \[ \large \frac{a}{b}=0 \] since b cannot be zero, the only possibility is a=0.

OpenStudy (helder_edwin):

this means that your equation \[ \large \frac{x^2-2x+1}{x-2}=0 \] is just \[ \large x^2-2x+1=0 \]